Through (-4,1), parallel to x+3y=5
Answer by fcabanski(1391)

Parallel lines have the same slope. y=mx+b is the slope-intercept form of a line. m is the slope.


x+3y=5


3y = -x + 5
y = -x/3 + 5/3. The slope is -1/3.


The slope point form of a line equation is y - y(1) = m(x-x(1)) where m is the slope and (x(1),y(1)) is a point on the line. Plug in the info.


y-1 = -1/3(x-(-4)


y-1 = -x/3 -4/3


y = -(1/3)x -1/3